# Deceptive Sequential Decision-Making via Regularized Policy Optimization

### v2

@@ IV-A3 Equivocal Deception
From Definition 7, a policy that is equivocal deceptive seeks to lead an adversary to a belief that two states are equally likely to be goal states. In this example, we seek to lead an adversary to infer that Agent0 and Agent1 are equally important. To achieve this, we aim to manipulate the occupancy measure $\{x(s,a)\}_{s\in\mathcal{S},a\in\mathcal{A}}$ to influence the metric $L^{i}$ so that $L^{i}$ of Agent0 and Agent1 have similar values.
The third row of Figure 2 shows that this goal is achieved with larger values of $\beta$. With $\beta\geq 300$, the likelihoods of Agent0 and Agent1 are closely aligned under all three types of IRL, while the other two decoy agents remain unimportant. Notably, the likelihoods of Agent0 and Agent1 estimated by Deep IRL are indistinguishable with $\beta\geq 300$. This setup provides equivocal deception by deceiving an adversary into believing there is no difference in importance between Agent0 and Agent1.
Figure 2: The adversary’s estimate of the real agents under three types of deception and with three types of IRL used by the adversary. Each plot shows results averaged over 5050 runs. When $\beta=0$, no deception is implemented, and the adversary correctly infers that the real agent is the most important. In diversionary deception, with $\beta\geq 0.3$, the adversary’s inference of the most important agent will be one of the decoy agents, confirming that the policy satisfies the definition of diversionary deception. In targeted deception, with $\beta\geq 0.2$ the likelihood of the target decoy agent is greater than that of the real agent, implying that deception misleads the adversary into believing that the target decoy agent is the real one. Thus, the policy satisfies the definition of targeted deception. In equivocal deception, with $\beta\geq 300$ the likelihoods of the real agent and the chosen decoy agent remain close to each other, which satisfies the definition of equivocal deception.
@@ IV-B Performance
In this section, we evaluate the performance of the system when deception is implemented with a policy $\pi_{d}$ obtained from Optimization Problems 45, or 7. We compare the revenue under deception with the optimal revenue. We also empirically verify the bounds on the revenue loss given in Theorems 12, and 3.
As shown in Figure 3, all three types of deception can be achieved with only small losses in revenue. In diversionary deception, with the choice of $\beta=0.3$, we still recover $97.39\%$ of the optimal revenue, which is less than a $3\%$ reduction compared to the revenue without deception. Similarly, in targeted deception with $\beta=0.2$ we find that the deceptive policy recovers $98.69\%$ of the optimal revenue, and thus the loss in revenue in this case is less than $2\%$ compared to the revenue without deception. In equivocal deception, with the choice of $\beta=300$ we observe that the deceptive policy attains $98.30\%$ of the optimal revenue, corresponding to a revenue loss of less than $2\%$ compared to the revenue without deception. Under each form of deception, we find that at values of $\beta$ that generate deceptive policies, the resulting policies incur minimal losses in revenue, highlighting that we can achieve both deception and high performance simultaneously.
Figure 3: The revenue obtained from deceptive policies under diversionary, targeted, and equivocal deception. In diversionary deception, the system recovers $97.39\%$ of the optimal revenue with $\beta=0.3$. With targeted deception and $\beta=0.2$, the system recovers $98.69\%$ of the optimal revenue. For equivocal deception with $\beta=300$, the system recovers $98.30\%$ of the optimal revenue. In all types of deception, less than $3\%$ of revenue is lost, which outperforms worst-case bounds on performance loss.
@@ V Conclusion
This study demonstrates that active deception strategies can effectively mislead adversaries in fully observable systems with formal performance guarantees under a specified degree of deception. We have developed analytic bounds for performance losses in terms of total accumulated rewards and verified the performance of deceptive decision policies in simulations. By ensuring reliable operation under deception, our findings enhance the practical applicability of deception in securing critical systems. Future work will explore tradeoffs between optimality and deception under a broader range of adversarial models and alternate definitions of deception.
@@ Appendix A Proposition 1
$\gamma\sum_{j\in\mathcal{S}}\sum_{s\in\mathcal{S}}\sum_{a\in\mathcal{A}}\mathcal{T}(s,a,j)x(s,a)$ $=\gamma\sum_{j\in\mathcal{S}}\sum_{s\in\mathcal{S}}\sum_{a\in\mathcal{A}}\mathcal{T}(s,a,j)\sum_{k\in\mathcal{S}}\alpha(k)\sum_{n=1}^{\infty}\gamma^{n-1}P^{\pi}_{n,k}(s,a)$ (26)
@@ Appendix B Proof of Theorem 1
$\sum_{s\in\mathcal{S}}\sum_{a\in\mathcal{A}}\Big{[}r(s,a)(x_{d}(s,a)-x^{*}(s,a))+\beta\big{(}x_{d}(s,a)-x^{*}(s,a)\big{)}^{2}\Big{]}\geq 0.$ (34)
@@ Appendix C Proof of Theorem 2
$L_{\pi_{d}}\leq\frac{\beta}{R^{*}}\left[\sum_{s\in\mathcal{S}}\sum_{a\in\mathcal{A}}\Big{(}x^{*}(s,a)^{2}\right.-2x_{tar}(s,a)x^{*}(s,a)\Big{)}\left.-\frac{(1-\gamma)^{-2}}{|\mathcal{S}||\mathcal{A}|}+2(1-\gamma)^{-1}\max_{s\in\mathcal{S},a\in\mathcal{A}}x_{tar}(s,a)\right],$ (42)

### v3

@@ IV-A3 Equivocal Deception
From Definition 7, a policy that is equivocal deceptive seeks to lead an adversary to a belief that two states are equally likely to be goal states. In the context of MTD, we seek to lead an adversary to infer that the real agent(s) and the target decoy agent(s) are equally important. To verify this goal, we consider 2 scenarios with different ratios of real and decoy agents, where we have one real and one target decoy agents in scenario 1, and one real and two target decoy agents in scenario 2. To achieve equivocal deception, we aim to manipulate the occupancy measure $\{x(s,a)\}_{s\in\mathcal{S},a\in\mathcal{A}}$ to influence the metric $L^{i}$ so that $L^{i}$ of the real and target decoy agents have similar values.
Figure 4: The adversary’s estimate of $L^{i}$ of each agent under equivocal deception with three types of IRL. Each plot shows results averaged over $100$ runs. We used $\textit{number of trajectories}=500$ and $\textit{length of trajectories}=500$ for all simulations, $\textit{epoch}=20$, $\textit{learning rate}=0.5$ for MaxEnt IRL, and $\textit{epoch}=50$, $\textit{learning rate}=0.1$ and two fully connected hidden layers with dimension $64$ and $32$ for Deep IRL. The blue line represents the real agent, the red line(s) represents the target decoy agent(s), and the gray lines represent decoy agents. When $\beta=0$, no deception is implemented, and the adversary correctly infers that the real agent is the most important. With $\beta\geq 50$ the likelihoods of the real agent and the target decoy agent(s) remain close to each other, which satisfies the definition of equivocal deception.
Deception Parameter $\beta$ 0 50 100 150 200 250 300 350 400 450 500
Scenario 1 Real 99 36 44 36 26 19 50 56 58 45 45
Target decoy 0 64 56 64 74 81 50 44 42 55 55
Decoy 1 0 0 0 0 0 0 0 0 0 0 0
Decoy 2 0 0 0 0 0 0 0 0 0 0 0
Decoy 3 1 0 0 0 0 0 0 0 0 0 0
Adversary’s Utility 99 -28 -12 -28 -48 -62 0 12 16 -10 -10
Scenario 2 Real 100 34 38 18 10 13 27 12 15 6 6
Target decoy 1 0 26 25 41 47 36 19 28 41 33 32
Target decoy 2 0 40 37 41 43 51 54 60 44 61 62
Decoy 1 0 0 0 0 0 0 0 0 0 0 0
Decoy 2 0 0 0 0 0 0 0 0 0 0 0
Adversary’s Utility 100 -32 -24 -64 -80 -74 -46 -76 -70 -88 -88
TABLE IV: Adversary’s estimation of the real agent and total utility when using Deep IRL under equivocal deception
Figure 4 shows that equivocal deception is achieved with larger values of $\beta$. While Apprenticeship learning and MaxEnt IRL shows some fluctuation in their estimation, in general, with $\beta\geq 50$, the likelihoods of the real agent and all of the target decoy agent(s) are closely aligned, while the other decoy agents remain unimportant. Table IV shows that the adversary attacks the real agent less often with $\beta\geq 50$ compared to the case with no deception, resulting in a decrease in the adversary’s utility. This setup provides equivocal deception by deceiving an adversary into believing there is no difference in importance between the real and the target decoy agents.
@@ IV-B Performance
In this section, we evaluate the performance of the system when deception is implemented with a policy $\pi_{d}$ obtained from Optimization Problem 45, or 7. We quantify the revenue lost by introducing each type of deception and verify the bounds on the revenue loss given in Theorems 12, and 3.
As shown in Figure 5, all three types of deception can be achieved with only small losses in revenue. In diversionary deception, with the choice of $\beta=0.4$, we still recover $98.04\%$ and $99.07\%$ of the optimal revenue, which is less than a $2\%$ reduction compared to the revenue without deception. Similarly, in targeted deception with $\beta=0.02$ we find that the deceptive policy recovers $99.95\%$ and $99.90\%$ of the optimal revenue, and thus the loss in revenue in this case is less than $1\%$ compared to the revenue without deception. In equivocal deception, with the choice of $\beta=50$ we observe that the deceptive policy attains $99.99\%$ and $99.98\%$ of the optimal revenue, corresponding to a revenue loss of less than $1\%$ compared to the revenue without deception. Under each form of deception, we find that at values of $\beta$ that generate deceptive policies, the resulting policies incur minimal losses in revenue, highlighting that we can achieve both deception and high performance simultaneously.
Figure 5: The revenue obtained from deceptive policies under diversionary, targeted, and equivocal deception. In diversionary deception, the system recovers $98.04\%$ and $99.07\%$ of the optimal revenue with $\beta=0.4$. For targeted deception with $\beta=0.02$, the system recovers $99.95\%$ and $99.90\%$ of the optimal revenue. For equivocal deception with $\beta=50$, the system recovers $99.99\%$ and $99.98\%$ of the optimal revenue. In all types of deception, less than $2\%$ of revenue is lost, which outperforms worst-case bounds on performance loss, verifying Theorem 12 and 3.
For practical implementation, Theorems 12 and 3 can provide a guide for selecting parameter $\beta$. For some scenarios where we have a maximum allowable performance loss, we can find a range of $\beta$ that attains required performance leveraging Theorems 12 and 3. In this example, the performance loss in scenario 2 of each types of deception are bounded by
$L_{\pi_{div}}\leq 0.82432\frac{\beta}{R^{*}},\quad L_{\pi_{tar}}\leq 0.62866\frac{\beta}{R^{*}},\quad L_{\pi_{equ}}\leq 0.00042\frac{\beta}{R^{*}}.$ (23)
For example, to achieve $80\%$ of it’s optimal performance, we need $0\leq\beta\lesssim 0.243$ for diversionary deception, $0\leq\beta\lesssim 0.168$ for targeted deception, and $0\leq\beta\lesssim 239.7$ for equivocal deception. This indicates that the choice of $\beta\simeq 0.243$$\beta\simeq 0.168$ and $\beta\simeq 239.7$ can provide the most deceptive policy under allowable performance loss.
From Table IIIII, and IV, we find $\beta=0.2$ in diversionary deception (scenario 2) drops the adversary’s utility from $100$ to $34$, $\beta=0.16$ in targeted deception (scenario 2) drops it to 100-100, and $\beta=200$ in equivocal deception (scenario 2) drops it to 80-80, while $80\%$ of its optimal performance is theoretically guaranteed in all cases. However, in practice, Figure 5 shows that the actual performance outperforms these bounds on performance loss in all three types of deception.
@@ V Conclusion
This study demonstrates that active deception strategies can effectively mislead adversaries in fully observable systems with formal performance guarantees under a specified degree of deception. We have developed analytic bounds for performance losses in terms of total accumulated rewards and verified the performance of deceptive decision policies with various scenarios in simulations. By ensuring reliable operation under deception, our findings enhance the practical applicability of deception in securing critical systems. Future work will explore tradeoffs between optimality and deception under a broader range of adversarial models, including some model that can expect and react to deception, and alternate definitions of deception.
Acknowledgments
Disclosure Statement
The authors report there are no competing interests to declare.
Data Availability Statement
The data that support the findings of this study were generated by the authors and are reproducible using the code publicly available at https://anonymous.4open.science/r/DeceptionMTD-2980. No external datasets were used.
@@ Appendix A Proposition 1
$\gamma\sum_{j\in\mathcal{S}}\sum_{s\in\mathcal{S}}\sum_{a\in\mathcal{A}}\mathcal{T}(s,a,j)x(s,a)$ (27)
$=$ $\gamma\sum_{j\in\mathcal{S}}\sum_{s\in\mathcal{S}}\sum_{a\in\mathcal{A}}\mathcal{T}(s,a,j)\sum_{k\in\mathcal{S}}\alpha(k)\sum_{n=1}^{\infty}\gamma^{n-1}P^{\pi}_{n,k}(s,a)$
@@ Appendix B Proof of Theorem 1
$\sum_{s\in\mathcal{S}}\sum_{a\in\mathcal{A}}\Big[r(s,a)(x_{d}(s,a)-x^{*}(s,a))\beta\big(x_{d}(s,a)-x^{*}(s,a)\big)^{2}\Big]\geq 0.$ (35)